OpenStudy (anonymous):

Simplify square root 72 minus 3 times square root 12 plus square root of 192 ???

OpenStudy (tkhunny):

\(\sqrt{72} = \sqrt{2\cdot 36} = \sqrt{2}\cdot\sqrt{36} = \sqrt{2}\cdot(6) = 6\sqrt{2}\) You do 12 and 192 and magical things should pop out.